OpenStudy (hannah_ryder):

The four blades of a helicopter meet at right angles and are all the same length. The distance the tips of two adjacent blades is 36 feet. How long is each blade? Round your answer to the nearest tenth.

OpenStudy (hannah_ryder):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Apply Pythagorean theorem: Let x = length (feet) of a blade then \[x^2 + x^2 = 36^2\]

OpenStudy (shadowlegendx):



OpenStudy (kendricklamar2014):

Substitute to get: \[2x^2 = 36^2\] Now solve..

OpenStudy (hannah_ryder):

x = 25.5 feet @KendrickLamar2014

OpenStudy (kendricklamar2014):

That is correct :)

OpenStudy (hannah_ryder):

Thank you ^_^

OpenStudy (kendricklamar2014):

Your welcome

OpenStudy (hannah_ryder):

Thanks for your help as well @ShadowLegendX